OpenStudy (anonymous):

Harry had $32. He spent all the money on buying 3 notebooks for $x each and 4 packs of index cards for $y each. If Harry had bought 5 notebooks and 5 packs of index cards, he would have run short of $18.

OpenStudy (anonymous):

A student concluded that the price of each notebook is $5 and the price of each pack of index cards is $1. Which statement best justifies whether the student's conclusion is correct or incorrect? The student's conclusion is incorrect because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 50 is (8, 2). The student's conclusion is incorrect because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 18 is (8, 2). The student's conclusion is correct because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 18 is (5, 1). The student's conclusion is correct because the solution to the system of equations 3x – 4y = 32 and 5x – 5y = 50 is (5, 1).

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@yoyogators

OpenStudy (campbell_st):

well you know that he spends his $32 this way 3x + 4y = 32 looking at the other option, you would been $50 to buy 5x + 5y = 50 so that will eliminate choices B and C so you task is to solve simultaneously 3x + 4y = 32 5x + 5y = 50 you could graph the lines, use the elimination method, the substitution method or simply substitute the points into the equations A has a solution x = 8 and y = 2 so substitute that into both equations and see if its true if not look at d x = 5, y = 1 substitute these values to see if its true.... hope it makes sense

OpenStudy (anonymous):

It does,l so it's A ? @campbell_st

OpenStudy (campbell_st):

that's correct, well done

OpenStudy (anonymous):

Thanks !